The Magazine Mass Marketing Company has received 1010 entries in its latest sweepstakes. They know that the probability of receiving a magazine subscription order with an entry form is 0.60.6. What is the probability that more than a fifth of the entry forms will include an order?
Answer #1

X ~ bin ( n , p)

Where n = 10 , p = 0.6

Binomial probability distribution is

P(X) = nCx * px * ( 1 - p)n-x

P(X > 5) = P(X >= 6)

= P(X = 6) + P(X = 7) + P(X = 8) + P(X = 9) + P(X = 10)

= 10C6 * 0.66 * ( 1 - 0.6)4 + 10C7 * 0.67 * ( 1 - 0.6)3 + 10C8 * 0.68 * ( 1 - 0.6)2 + 10C9 * 0.69 * ( 1 - 0.6) +

10C10 * 0.610 * ( 1 - 0.6)0

= 0.6331